40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 18 12:27:42 2019, gautam, Update, General, IFO status IMG_7343.JPGITMXOL.png
    Reply  Mon Mar 18 17:06:12 2019, gautam, Update, Optical Levers, ITMY HeNe replaced OLRIN.pdfOL_PIT.pdfOL_YAW.pdf
    Reply  Tue May 28 17:41:13 2019, gautam, Update, General, IFO status IMG_7444.JPG
Message ID: 14484     Entry time: Mon Mar 18 17:06:12 2019     In reply to: 14483
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: ITMY HeNe replaced 

Oplev HeNe was replaced this afternoon. We did some HeNe shuffling:

  1. A new HeNe was being used for the fiber illumination demo at EX. We took that out and decided to use it as the new ITMX HeNe. It had 2.6mW output at 632nm (measured with the Ophir power meter)
  2. Old ETMY HeNe was used for fiber illumination demo.
  3. Old ITMX HeNe was putting out no light - it will be disposed.

Attachment #1 shows the RIN and Attachment #2 and #3 show the PIT and YAW TFs with the new HeNe.

The ITMX Oplev path is still not great - the ingoing beam is within 2mm of clipping on a 2" lens used in the POX path, and there is a bunch of scattered red light everywhere. We should take the opportunity when the chamber is open to try and have a better layout (it may be tricky to optize without touching the two in-vacuum steering optics).

Quote:

I'll ask Chub to replace it this afternoon.

Attachment 1: OLRIN.pdf  65 kB  Uploaded Mon Mar 18 18:18:37 2019  | Hide | Hide all
OLRIN.pdf
Attachment 2: OL_PIT.pdf  11 kB  Uploaded Mon Mar 18 18:18:47 2019  | Hide | Hide all
OL_PIT.pdf
Attachment 3: OL_YAW.pdf  12 kB  Uploaded Mon Mar 18 18:18:53 2019  | Hide | Hide all
OL_YAW.pdf
ELOG V3.1.3-